# Variational Bayesian latent estimation with compressive autoencoders

### v2

@@ 3.3 Discussion and position to related works
First, it is worth noting that this variational framework does not increase the restoration time compared to its deterministic counterpart MAP-z [19]. Indeed, after convergence of VBLE, sampling from the approximate posterior is fast since it only requires forward passes through the decoder, which can be parallelized within batches. Compared to other existing posterior sampling methods to solve inverse problems, MCMC techniques [33, 35] enable to sample from the true posterior instead of an approximation, but these methods require a substantial number of iterations, making them far slower than VBLE.
Concerning variational inference methods, they have been widely used to approximate inverse problem posterior distributions [43, 53]. But the approaches which combine variational inference with deep generative modelling are essentially designed in a (semi-)supervised manner to solve a specific inverse problem [53, 54, 55], while VBLE is able to solve different inverse problems with the same network. Then, latent optimization methods [19, 22, 23] yield, up to our knowledge, only punctual solutions, except from Holden et al. [35] which, however, is an MCMC method.
Regarding the use of CAEs as regularizers, the relationship between CAEs and VAEs have been established in the literature [38, 56]. But, up to our knowledge, this is the first time that CAEs are used as priors for image restoration tasks. Although CAEs are not as deep as state-of-the-art hierarchical VAEs and may not exhibit the same performance for image generation, we believe they make appropriate priors for image restoration. In particular, the parameter $\alpha$, which tunes the rate-distortion trade-off, provides additional flexibility and can be adjusted for a given inverse problem.
@@ 4 Numerical experiments
In this section, we evaluate the performance of VBLE using CAEs and compare the results with state-of-the-art baselines. We first detail the experimental setup, and then present the results on two datasets. The first one, FFHQ, is highly structured and thus well-suited for generative models, whereas the second one, BSD, contains natural images with more diversity. Regarding the baselines, we will focus on methods based on generative models for the former, and denoising networks for the latter. Finally, we assess the quality of our proposed approximate posterior compared to MCMC methods.
@@ 4.1 Experimental setup
Inverse problems. All experiments in this section are performed on subsets of BSD500 [41] and FFHQ [40] test datasets, both subsets composed of 47 images of size $256\times 256$. VBLE and the baselines are evaluated on three inverse problems: deblurring, single image super resolution (SISR), and inpainting. For deblurring, two Gaussian kernels of standard deviation $\sigma_{blur}\in\{1,3\}$ are tested as well as a motion blur kernel, for several Gaussian noise levels. We consider the noiseless case for SISR and inpainting with random masks, and a small amount of noise is added for inpainting with structured masks. SISR ×$\times 2$ and ×$\times 4$ is conducted with bicubic downsampling. We employ a $50\%$ random mask for inpainting experiments with random masks. For inpainting with structured masks, custom masks are designed on 5 test images of each dataset.
@@
Metrics. We use three metrics to measure the discrepancy between the restored and ground truth images: Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity [57] (SSIM) and Learned Image Patch Similarity [58] (LPIPS). The PSNR derives from the pixel-wise mean-squared error, while SSIM and LPIPS are respectively classical and deep learning-based perceptual metrics.
@@
Compressive autoencoder architectures. We employ pretrained networks from the compressAI [59] library at different bitrates. We choose Cheng et al. [47] model, which is a CAE with an hyperprior combining a second autoencoder with an autoregressive model, yielding state-of-the-art compression results. The networks from compressAI are pretrained on natural images. We finetune them for about 100k iterations on BSD and FFHQ train datasets.
@@
VBLE parameter settings. For VBLE using CAEs, the network with the most appropriate bitrate needs to be chosen for each inverse problem. We have trained models at 7 different bitrates. The bitrate parameter $\alpha$ to choose, as well as the regularization parameter $\lambda$ are tuned by a grid search algorithm on a defined 3 image validation set for each dataset. The other parameters are fixed, in particular, we use Adam optimizer with learning rate 0.1 during 1000 iterations for all restoration tasks. Lastly, the MMSE-x estimate is used for all comparisons as it yielded the best results.
Detailed experimental settings are given in the supplementary material, as well as extended FFHQ and BSD experiments and an ablation study about VBLE variants and hyperparameters.
@@ 4.2 FFHQ experiments
First, we demonstrate the relevance of VBLE using CAEs for restoring highly structured face images from FFHQ dataset. To assess the effectiveness of both the CAE architecture and VBLE algorithm, we provide restoration results with VBLE (denoted as VBLE wCAE in the experiments) and also with its deterministic counterpart MAP-z using the same CAE (denoted as MAP-z wCAE).
We compare our method to the latent optimization approach Bora and al. [19] (denoted as MAP-z wVAE), which seems a natural baseline for us, as well as the state-of-the-art diffusion-based method DiffPIR [18]. For DiffPIR, we use an available FFHQ pretrained diffusion model [60]. For Bora and al. method, we train two vanilla VAE models on FFHQ. VAE-V1 and V2 are two single latent variable VAEs without and with hyperprior, with a classical convolutional structure, with approximately the same number of parameters than the CAE and the same latent dimension. Although more elaborate VAEs are likely to perform better as long as their size do not hamper the latent optimization, here we aim to assess the benefit of the CAE architecture, with its specific hyperprior. Hence we choose to compare with VAEs having similar complexity. More details on these architectures are given in supplementary material.
FFHQ Deblur (motion) SISR ×$\times 4$ Inpainting (random)
Method PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
VBLE wCAE 32.08 0.2010 0.8751 31.26 0.2048 0.8699 36.98 0.0795 0.9623
MAP-z wCAE 32.01 0.2441 0.8717 31.08 0.2205 0.8681 36.91 0.0715 0.9618
MAP-z wVAE V1 27.95 0.4331 0.6582 30.24 0.2201 0.8499 35.21 0.0775 0.9443
MAP-z wVAE V2 29.44 0.3088 0.8112 29.53 0.2765 0.8226 31.01 0.2660 0.8513
DiffPIR 31.98 0.1793 0.8677 30.71 0.2042 0.8492 36.08 0.0653 0.9470
@@
Table 1: FFHQ results on diverse inverse problems: motion deblurring ($\sigma=7.65/255$), SISR ×$\times 4$, inpainting with random masks.
Figure 4: Qualitative image restoration results on FFHQ. From top to bottom: Gaussian deblurring ($\sigma_{blur}=1,\sigma=7.65/255$), SISR ×$\times 2$, inpainting.
@@
Quantitative results on FFHQ along with a visual comparison are provided in Tab. 1 and Fig. 4. First, the use of VBLE algorithm instead of a deterministic gradient descent not only enables posterior sampling but also enhances point estimation results, as VBLE with CAE yields better metrics than MAP-z with CAE and tends to exhibit sharper images. Then, Bora’s method MAP-z with VAEs are outperformed by VBLE, both quantitatively and qualitatively. VAE-V1 seems to lack structure while VAE-V2 lacks high frequencies, demonstrating the interest of the CAE structure and its hyperprior. Compared to DiffPIR, VBLE exhibits better PSNR and SSIM, and slightly lower LPIPS scores, which indicates that these methods have a slightly different position in the perception-distortion tradeoff [61]. Visually, both methods perform very well on FFHQ.

### v4

@@ 3.3 Discussion and position to related works
Regarding latent optimization methods, most of them consider point estimates [7, 34, 27]. Yet, several approaches have proposed posterior sampling or ensemble methods that enable uncertainty quantification. It may be through sampling the latent posterior using MCMC [37, 62] or, as in the case of PULSE [48], by running a spherical gradient descent multiple times with different random initializations. PULSE does not provide samples of the latent posterior, but several MAP estimates. On the contrary, [37, 62] sample from the latent posterior. In particular, in [37], the latent posterior is defined exactly as in VBLE, that is $p_{Z|Y}(z|y)\propto p(z)p_{Y|X}(y|D_{\theta}(z))$, thus the two methods are expected to produce similar posterior samples. Compared to those methods, VBLE is a Variational Inference (VI) method, that only approximates the posterior distribution, but with a single latent optimization of the variational parameters. This key difference leads to a significantly reduced computational load for VBLE compared to MCMC methods [37, 62] and PULSE [48], as drawing several samples requires several latent optimizations with PULSE, and typically more than $10^{6}$ iterations with MCMC algorithms. Furthermore, to transform the latent samples into images, we consider a deterministic decoder, that is $p_{\theta}(x|z)=\delta_{D_{\theta}(z)}(x)$ while [62] obtains image samples by sampling $p(x|z,y)\propto p(x|z)p(y|x)$. By doing so in VBLE, algorithm 1 would remain unchanged, but the decoding, in particular of Equations 13,14 would be modified: the images would be potentially a bit less blurry. However, our main concern is that the decoder distribution $p_{\theta}(x|z)=\mathcal{N}(x;D_{\theta}(z),\gamma^{2}I)$ classically used in VAEs and CAEs is not well adapted as the fixed variance is not informative enough. Alternative approaches to improve the decoder variance [28] could be used. As this is not straightforward to apply, this interesting perspective is left for future work.
Concerning VI methods, they have been widely used to approximate posterior distributions of inverse problem solutions [6, 65]. However, the approaches which combine VI with deep generative modeling are essentially designed in a (semi-)supervised manner to solve a specific inverse problem [65, 31, 33], while VBLE is able to solve different inverse problems with the same network. In [14], denoising is performed through score prior-guided deep VI, VI being used to approximate the noise model parameters. This requires optimizing the weights of a convolutional neural network (CNN) that predicts the noise parameters for each input image $y$ to be restored. In contrast, since VBLE assumes that the noise and the degradation are known, it only uses VI to approximate the latent posterior. For that purpose, the variational parameters are directly optimized using gradient descent. Hence, up to our knowledge, VBLE is the first latent optimization method which explicitly estimates the latent posterior with variational inference to solve general inverse problems. Moreover, the uniform posterior shape for VBLE with CAE, which mimics the encoder posterior, is novel and yields remarkable results for solving imaging inverse problems, as will be shown in the next section.
Regarding the use of CAEs as regularizers, the relationship between CAEs and VAEs has been established in the literature [3, 49]. But, as far as we know, this is the first time that CAEs are used as priors for image restoration tasks. Although CAEs are not as deep as state-of-the-art hierarchical VAEs and may not exhibit the same performance for image generation, we believe that they make appropriate priors for image restoration. In particular, the parameter $\alpha$, which tunes the rate-distortion trade-off, provides additional flexibility and can be adjusted for a given inverse problem.
@@ 4 Numerical experiments
In this section, we evaluate the performance of VBLE using CAEs. With these experiments, our purpose is twofold. First, we aim to show that VBLE, in particular with the use of CAEs, yields state-of-the-art results for image restoration tasks. Second, we wish to demonstrate the huge interest of the proposed method for posterior sampling, which can be crucial for many applications.
In that purpose, we first detail the experimental setup, and then present image restoration results on two datasets. The first one, FFHQ, is highly structured and thus well-suited for generative models, whereas the second one, BSD, contains natural images with more diversity. Then, we assess the quality of our proposed approximate posterior both quantitatively and qualitatively compared to state-of-the-art posterior sampling baselines. Finally, we conduct an ablation study about our method and its variants.
@@ 4.1 Experimental setup
All experiments in this section are performed on subsets of BSD500 [47] and FFHQ [41] test datasets, both subsets composed of 47 images of size $256\times 256$. VBLE and the baselines are evaluated on three inverse problems: deblurring, single image super-resolution (SISR), and inpainting. For deblurring, two Gaussian kernels of standard deviation $\sigma_{blur}\in\{1,3\}$ are tested as well as a motion blur kernel, for several Gaussian noise levels. We consider the noiseless case for SISR and inpainting with random masks, and a small amount of noise is added for inpainting with structured masks. SISR $\times 2$ and $\times 4$ is conducted with bicubic downsampling. We employ a $50\%$ random mask for inpainting experiments with random masks. For inpainting with structured masks, custom masks are designed on 5 test images of each dataset.
Metrics
@@
We use four metrics to measure the discrepancy between the restored and ground truth images: Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity [67] (SSIM), Learned Image Patch Similarity [72] (LPIPS) and Fréchet Inception Distance (FID) [35]. The PSNR derives from the pixel-wise mean-squared error, while SSIM and LPIPS are, respectively, classical and deep learning-based perceptual metrics. The FID is a no-reference perceptual metric. For each inverse problem, we evaluate the FID after extracting 49 patches of size $64\times 64$ from each restored image.
Compressive autoencoder architectures
@@
We employ pretrained networks from the compressAI [9] library at different bitrates. We choose two models, from Minnen et al. [50], denoted as mbt, and Cheng et al.[15], denoted as cheng. mbt is a CAE with a hyperprior combining a second autoencoder with an autoregressive model, but its encoder and decoder structures remain simple CNNs. cheng retains the same structure as mbt, but with more elaborate encoders and decoders as they possess self-attention layers, yielding state-of-the-art compression results. For mbt and cheng models, two structures, a light and a bigger one, exist [9], which are used respectively for low and high bitrates. For mbt, a high bitrate structure is used for $\alpha\geq 0.013$, with $\alpha$ the bitrate parameter, and a low bitrate structure for $\alpha\leq 0.0067$. For cheng, we have chosen the high bitrate structure for high and low bitrates, as it provides the best image restoration results. The number of parameters for each network is given in LABEL:tab:nparams. The networks from compressAI are pretrained on natural images. We finetune them for about 100k iterations on BSD and FFHQ train datasets.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
VBLE wCAE 34.48 0.9250 0.1473 38.03 29.57 0.8214 0.2859 74.12 32.22 0.8800 0.1933 40.70
MAP-z wCAE 34.17 0.9152 0.1684 48.88 29.44 0.8176 0.2977 92.65 32.01 0.8717 0.2441 64.08
VBLE wCAE (mbt) 34.21 0.9177 0.1600 43.44 28.70 0.7686 0.3212 102.96 31.97 0.8741 0.2149 48.15
MAP-z wVAE V1 32.78 0.8868 0.1990 71.43 28.72 0.7961 0.3242 120.15 27.95 0.6582 0.4331 192.81
MAP-z wVAE V2 30.91 0.8486 0.2688 66.01 28.47 0.7849 0.3263 79.12 29.44 0.8112 0.3088 73.55
PULSE 29.99 0.8289 0.2848 64.42 27.83 0.7670 0.3429 81.00 28.36 0.7848 0.3266 74.85
DiffPIR 33.92 0.9049 0.1368 36.12 29.41 0.8102 0.2691 100.98 31.98 0.8677 0.1793 42.50
DDRM 32.22 0.8831 0.1785 33.84 24.74 0.7051 0.2931 41.01 x x x x
SISR and Inp. SISR $\times 2$ SISR $\times 4$ Inpainting (random)
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
VBLE wCAE 36.16 0.9468 0.1251 40.44 31.18 0.8670 0.2290 56.21 35.94 0.9584 0.0638 31.01
MAP-z wCAE 36.30 0.9516 0.0995 34.59 31.08 0.8681 0.2205 62.91 36.91 0.9618 0.0715 30.42
VBLE wCAE (mbt) 35.99 0.9514 0.0956 38.64 31.01 0.8680 0.2258 60.95 37.04 0.9643 0.0738 33.73
MAP-z wVAE V1 34.35 0.9311 0.0893 38.51 30.24 0.8499 0.2201 75.74 35.21 0.9443 0.0775 35.00
MAP-z wVAE V2 31.20 0.8534 0.2592 64.24 29.53 0.8226 0.2765 66.55 31.01 0.8513 0.2660 66.29
PULSE 29.79 0.8203 0.2989 66.85 27.11 0.7395 0.3927 85.81 30.06 0.8300 0.2858 64.37
DiffPIR 34.94 0.9269 0.1089 34.63 30.71 0.8492 0.2042 68.34 36.08 0.9470 0.0653 30.93
DDRM 35.81 0.9479 0.0640 27.19 30.80 0.8608 0.1749 38.65 34.43 0.9372 0.0814 24.54
Table 1: FFHQ results on diverse inverse problems. For methods using CAEs, (mbt) means that mbt CAE [50] is used, otherwise cheng CAE [15] is used. $\sigma$ is the noise level used for deblurring. $\sigma_{k}$ is the kernel standard deviation for Gaussian deblurring.
VBLE parameters
@@
For VBLE using CAEs, the network with the most appropriate bitrate needs to be chosen for each inverse problem. We have trained models at 7 different bitrates. The bitrate parameter $\alpha$ to choose, as well as the regularization parameter $\lambda$, are tuned by a grid search algorithm on a 3-image validation set for each dataset. The other parameters are fixed, in particular, we use Adam optimizer with a learning rate of 0.1 during 1000 iterations for all restoration tasks. Lastly, the MMSE-x estimate, computed by averaging 100 VBLE posterior samples, is used for all comparisons as it yielded the best results.
Image restoration parameters
To tune image restoration hyperparameters of VBLE and of the baselines, a grid search algorithm has been used on separate 3 image sets of FFHQ and BSD test datasets. The values of all the VBLE parameters for the different problems are given in the appendix, as well as the parameters that have been tuned or fixed for each baseline.
@@ 4.2 FFHQ experiments
Figure 4: Visual results on FFHQ. From top to bottom: : motion deblurring ($\sigma=7.65/255$), SISR $\times 4$, inpainting with random masks.
BSD - $\sigma=7.65$ Deblur (Gauss., $\sigma_{k}=1$) Deblur (Gauss., $\sigma_{k}=3$) Deblur (Motion)
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
VBLE wCAE 29.62 0.8598 0.1901 61.38 24.66 0.6433 0.3956 89.22 28.38 0.8158 0.2356 64.91
MAPz wCAE 29.28 0.8405 0.2215 68.03 24.48 0.6347 0.4193 119.67 28.07 0.7896 0.3111 101.16
PnP-ULA 27.71 0.7809 0.2675 67.50 23.75 0.5913 0.4592 114.23 27.17 0.7402 0.2643 71.56
DPIR 29.72 0.8623 0.2127 73.76 24.41 0.6427 0.4315 134.92 28.69 0.8285 0.2606 83.44
PnP-ADMM 28.96 0.8456 0.2155 67.93 24.24 0.6333 0.4350 131.58 28.16 0.8045 0.3349 127.81
DiffPIR 29.26 0.8420 0.1832 55.34 24.58 0.6405 0.3605 98.93 28.15 0.7955 0.2230 61.19
DDRM 27.03 0.7763 0.2238 56.12 20.32 0.4747 0.4376 71.75 x x x x
BSD SISR $\times$2 SISR $\times$4 Inpainting (Random)
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
VBLE wCAE 29.55 0.8748 0.1802 65.72 25.53 0.7032 0.3507 99.09 30.21 0.9070 0.1207 56.60
MAPz wCAE 29.65 0.8757 0.1675 62.53 25.38 0.6966 0.3462 100.87 30.45 0.9091 0.1247 56.25
PnP-ULA 28.42 0.8763 0.1659 57.40 24.82 0.6894 0.3730 108.86 26.29 0.8122 0.2497 81.46
DPIR 29.55 0.8784 0.1706 61.35 25.18 0.6937 0.3652 114.41 31.58 0.9316 0.0659 58.20
PnP-ADMM 29.31 0.8678 0.1907 66.94 24.77 0.6659 0.4227 139.96 31.69 0.9322 0.0675 53.68
DiffPIR 29.20 0.8539 0.1645 57.78 25.03 0.6771 0.3164 88.83 30.41 0.9018 0.0919 55.94
DDRM 28.73 0.8691 0.1284 51.27 25.12 0.6895 0.2885 64.57 29.48 0.8754 0.1268 52.59
Table 2: BSD results on diverse inverse problems. $\sigma$ is the noise level used for deblurring. $\sigma_{k}$ is the kernel standard deviation for Gaussian deblurring.
First, we demonstrate the relevance of VBLE using CAEs for restoring highly structured face images from FFHQ dataset. To assess the effectiveness of both the CAE architecture and VBLE algorithm, we provide restoration results with VBLE (denoted as VBLE wCAE in the experiments) et also with its deterministic counterpart MAP-z using the same CAE (denoted as MAP-z wCAE). Results are provided with cheng and mbt CAE structures.
FFHQ Baselines
We compare our method to several latent optimization approaches. First, MAP-z wVAE, proposed by Bora et al. [7] is the deterministic counterpart of VBLE, which seems a natural baseline. We also compare to PULSE [48], which performs spherical gradient descent in the latent space. Then, we compare to the state-of-the-art diffusion-based methods DiffPIR [73] and DDRM [42]. We use 100 Neural Function Evaluations (NFEs) for both methods. DiffPIR’s and DDRM’s outputs are a sample from the posterior distribution $p(x|y)$. Hence, they can be considered as posterior sampling methods, even if sampling from the posterior requires running the restoration process multiple times. When evaluating image restoration performance, we choose to run DiffPIR and DDRM only once, as is classically done in the literature. For DDRM, we do not perform motion blur as it is not implemented in the source code. For these baselines, we use an available FFHQ pre-trained diffusion model [16]. For Bora et al. method [7], we train two vanilla VAE models on FFHQ, VAE-V1, and VAE-V2. With these models, our aim is to demonstrate the effectiveness of the two latent variable structure specific to CAEs. Hence, these models are vanilla VAEs, that is VAEs with a single latent variable. Their structure has been designed so that it has the same depth, same number of hidden channels and same latent dimension as mbt with (respectively without) hyperprior in the case of VAE-V1 (respectively VAE-V2). Hence, VAE-V1 encoder possesses 5 convolutional layers and VAE-V2 7 convolutional layers. The latent dimensions of VAE-V1 and V2 correspond to $z$ (first latent variable) and $h$ (second latent variable) dimensions of mbt CAE. Both models have been trained on FFHQ using a learning rate of $1e^{-4}$ during at least 300k iterations. We choose a Gaussian assumption for the decoder, that is $p_{\theta}(x|z)=\mathcal{N}(D_{\theta}(z),\gamma^{2}I).$ The variance $\gamma^{2}$ is considered a global parameter and is learned with the other network parameters. VAE-V2 model is used for PULSE [48] baseline. Although more elaborate VAEs are likely to perform better as long as their size does not hamper the latent optimization, here we aim to assess the benefit of the CAE architecture, with its specific hyperprior. Hence we choose to compare with VAEs having similar complexity.
Results
@@
Quantitative results on FFHQ are provided in Tab. 1. First, VBLE exhibits state-of-the-art results compared to the baselines. Compared to DiffPIR, VBLE exhibits better PSNR and SSIM, a similar FID, and slightly lower LPIPS scores, which indicates that these methods have a slightly different position in the perception-distortion tradeoff [5]. DDRM outperforms all methods in FID, but is below VBLE and DiffPIR for the other metrics. Then Bora’s method MAP-z with VAEs and PULSE are outperformed by VBLE. Furthermore, VBLE with CAE yields better metrics than MAP-z with CAE, hence the use of VBLE algorithm instead of a deterministic gradient descent not only enables posterior sampling but also improves point estimation results. Additionally, visual comparisons are provided in Fig. 4. VBLE, DDRM and DiffPIR perform very well on FFHQ. Compared to VBLE, the diffusion-based methods yield particularly good reconstruction on details such as eyes and teeth. This can be partly explained by the MMSE-x estimate chosen for VBLE. Indeed, it corresponds to the posterior mean, which tends to be smoother but more faithful to the measurement. VAE-V1, which imitates a CAE without hyperprior, seems to poorly restore structure while VAE-V2, which has the same depth as a CAE with hyperprior but has a single latent variable, lacks high frequencies. This demonstrates the interest of the CAE structure and its hyperprior: its two-latent variable structure achieves retrieving both semantic information and finer details.